a varible y varies directly with the square of x. if x=2 when y-4, find the constant of proportionality, k.
how do you find that out? I am confused. please help. thanks so much!

Found 2 solutions by stanbon, tutorcecilia:
Answer by stanbon(48568)

a varible y varies directly with the square of x. if x=2 when y=4, find the constant of proportionality, k.

Based on the 1st sentence y=kx^2
You need to find "k".
Based on the 2nd sentence 4=k(2^2)
k=4/4=1
Comment: I noticed you might have meant to type y=-4
If so, you would have -4=k(2^2).
Then k would equal -1
Cheers,
Stan H.

Answer by tutorcecilia(2144)

Since this is direct variation, use the formula:
y=x^2k ["k" is the constant that does not change]
This means that as y changes, x changes. If y is increased, x is increased. If y decreases, x will also decrease. "k" will stay the same.
.
Since x=2 and y=4, than plug-in the values and solve for "k"
4=2^2k
4=4k
4/4=4k/4
1=k
So, y=x^(1)
.
For an indirect variation, use the formula:
y=k/x [And plug-in the values]
This means that as y changes, x changes inversely. That is, as y increases, x decreases. If y decreases, x increases. In either case, "k" will stay the same.